OpenStudy (anonymous):

solve the equation sqrt(10-x) + x = 8

OpenStudy (jamesj):

Write this as sqrt(10-x) = 8 -x and square both sides

OpenStudy (anonymous):

Solve: \[\sqrt{10-x} + x = 8\]

OpenStudy (anonymous):

(\[(\sqrt{10-x} - x = 8 )^{2}\] (\[(\sqrt{10-x}) = (8-x)^2 10 - x = 64 - 16x +x^2 10-x = x^2 - 16x +64 10-x=(x-8)^2

OpenStudy (anonymous):

thats what i got so far! but i got stuck

OpenStudy (jamesj):

Now multiply out the (x-8)^2 and then collect terms to form a quadratic equation

OpenStudy (anonymous):

10-x = x^2 - 16x +64 ??

OpenStudy (anonymous):

x=6

OpenStudy (jamesj):

x = 6 is one root, yes. Is there another?

OpenStudy (anonymous):

oh no that was it. thanks !